DYNAMICS OF LINEAR AND T-SHAPED ArI_{2} DISSOCIATION UPON B←X EXCITATION: A DISPERSED INTRACAVITY LASER-INDUCED FLUORESCENCE STUDY

\underline{AMY E. STEVENS MILLER}, CHENG-CHI CHUANG, KELLY J. HIGGINS, HENRY FU and WILLIAM KLEMPERER, Department of Chemistry, Harvard University, Cambridge, MA 02138.

We have recorded the dispersed fluorescence spectra of the linear and the previously well-studied T-shaped isomeric ArI_{2} complexes following B←X excitation. The dispersed fluorescence features originate from the free I_{2} B→X emission and reveal the I_{2} product vibrational state distribution. The vibrational distribution of photofragment I_{2}(B) produced from dissociation of the T-shaped isomer shows a sharp Poisson distribution, beginning with Δv′ = −3, which is nearly independent of excitation energy. The vibrational distribution of I_{2}(B) from the linear isomer is quite different. For excitation energies greater than that corresponding to I_{2} (B,v’=22), the product distribution is quite similar to that observed for excitation above the dissociation energy of the B state, i.e., it shows a broad, nearly gaussian distribution of I_{2}(B) vibrational states, well below the excitation energy. For excitation energies below that, the distribution becomes increasingly like the Poisson distribution observed for the T-shaped isomer. Bond dissociation energies for linear and T-shaped isomers are determined for the spectra. The dissociation time scale for the two isomers differs by about 3 orders of magnitude. To model the product state distribution of the linear isomer and therefore gain further insight into the Ar-I_{2} interaction potential, a simple trajectory calculation based upon the classical dynamics of a compressed oscillator (I−I) ejecting a freely attached particle (Ar) is in progress. Results from the calculation are in good accord with the experiment.